OpenStudy (anonymous):

Find where the function is concave up: f(x) = 3x4 - 2x3 – 3x

OpenStudy (anonymous):

So far I did find f '(x)which equals 12x^3-6x^2-3

OpenStudy (amistre64):

f'' is a good test for cave up....which is a MIN find your "zeros" for the f' and that will tell you what "x" values you need to use for the next part..

OpenStudy (amistre64):

3(4x^3 -2x^2 -1) narrows it down. 4x^3 -2x^2 -1 = 0 x^2(4x-2) = 1 might help

OpenStudy (amistre64):

lets go ahead and find the f'' might be easier

OpenStudy (amistre64):

f'' = 36x^2 - 12x now we can partition this out to determine what values we get for the number line

OpenStudy (amistre64):

12x(3x -1).... we have zero values at 0 and 1/3 these are out inflections if were lucky <........0.........1/3..........> - + + - - + ---------------------- + - + getting warmer

OpenStudy (amistre64):

0 and 1/3 are our inflections were the concavity changes (-inf,0) = cave up (0,1/3) = cave down (1/3, inf) = cave up so this thing looks like a "W"